OpenStudy (anonymous):

Solve for x . -3=4(x-3)-7x Simplify your answer as much as possible.

OpenStudy (anonymous):

-3=4(x-3)-7x -3=4x-12-7x -3=-3x-12 +12 +12 9=-3x /-3 x=-3